# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 1.4. Organization
Part I of this paper presents background and recalls the work of Culver–Quigley in [CQ21]. In Section 2, we recall relevant facts about the stable motivic homotopy category $\text{SH}(F)$, the kq-resolution, and motivic Brown–Gitler comodules. In Section 3, we prove a novel result about symplectic K-theory. In Section 4, we outline a process for computing the ring of cooperations $\pi_{**}^{F}(\text{kq}\otimes\text{kq})$. Much of the material of these section is contained [CQ21, Section 3], but we phrase things in slightly more generality, which will be useful to a broader audience. In Section 5, we review the computation of $\pi_{**}^{\mathbb{C}}(\text{kq}\otimes\text{kq})$ as a warm up for the main contents of our work.
Part II of this paper contains the computation of the ring of cooperations $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$ and applications. In Section 6, we give describe the ring $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ and some modules which will frequently appear. In Section 7, we compute the ring $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ by a series of algebraic Atiyah–Hirzebruch spectral sequences. In Section 8, we compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$ and apply our results to describe the $E_{1}$-page of the kq-resolution.
@@ 1.5. Notation and Conventions
  • KQ denotes the motivic spectrum representing Hermitian K-theory, kq denotes its very effective cover, and ksp denotes the very effective cover of $\Sigma^{4,2}\text{KQ}$
  • $\text{Ext}_{B}(\mathbb{M}_{2}^{F},X)$ will be abbreviated by $\text{Ext}_{B}(X)$.
  • We grade Ext groups as $(s,f,w)$, where $s$ is the stem, $f$ is the Adams filtration, and $w$ is the motivic weight. We also let $cw=s-w$ denote the coweight, sometimes referred to in the literature as the Milnor-Witt degree.
@@
[…]
@@ 1.6. Acknowledgments
The work presented here constitutes a part of the author’s forthcoming thesis. The author thanks their advisors, Kyle Ormsby and John Palmieri, for their indispensable wisdom and valuable advice. The author also thanks J.D. Quigley for his generosity in sharing insight into this problem, and thanks Sarah Petersen and Alex Waugh for helpful discussions throughout this project. Finally, the author thanks Noah Wisdom for providing valuable feedback on an early draft.
@@ 2.1. The kq-resolution
It was shown in [Hor05] that there is an $(8,4)$-periodic motivic ring spectrum $\textup{KQ}\in\text{SH}(F)$ representing Hermitian K-theory, in that $\pi^{F}_{s,w}\textup{KQ}=\textup{K}^{\text{h}}_{s-2w}(F)$. In particular, $\pi_{8n,4n}^{F}\textup{KQ}=\textup{GW}(F)$, the Grothendieck–Witt ring of quadratic forms over $F$, and the unit map $\mathbb{S}\to\text{KQ}$ induces Morel’s isomorphism [RSØ19]
$\pi^{F}_{-n,-n}\mathbb{S}\cong\textup{K}^{\textup{MW}}_{n}(F).$
Let kq denote the very effective cover of KQ [ARØ20] (see Section 3 for more details on the very effective filtration). There is a canonical Adams tower associated to the unit map $\mathbb{S}\to\textup{kq}$:
@@
Applying $\pi^{F}_{**}(-)$ yields the motivic Adams spectral sequence kq-$\textbf{mASS}(\mathbb{S})$ known as the kq-resolution. Its properties were first studied by Culver–Quigley [CQ21]. We remark that we are working in the 2-complete category.
Theorem 2.1.
[CQ21, Thm 2.1] The kq-resolution is a strongly convergent spectral sequence of the form
$E_{1}^{s,f,w}=\pi^{F}_{s+f,w}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes f})\implies\pi^{F}_{s,w}\mathbb{S}_{2}^{\wedge}.$
@@
We can also look at the kq-resolution in the integral, uncompleted setting. The unit map $\mathbb{S}\to\textup{kq}$ gives an isomorphism $\Pi_{0}^{F}(\textup{kq})\cong\Pi_{0}^{F}(\mathbb{S})$, where $\Pi_{n}^{F}(X):=\bigoplus_{k\in\mathbb{Z}}\pi_{n+k,k}^{F}X$ denotes the $k$-th coweight stem (also known as the $k$-th Milnor–Witt stem). This implies that $\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}$ is $n$-connected in Morel’s homotopy $t$-structure. In fact, this is enough to show that $\mathbb{S}_{\textup{kq}}^{\wedge}\simeq\mathbb{S}$.
[…]
Let bo denote the connective cover of the classical real K-theory spectrum KO. The kq-resolution is the motivic analogue of the bo-resolution in the following way. The $E_{1}$-page of the bo-resolution is of the form
$E_{1}^{s,f}=\pi_{s+f}(\textup{bo}\otimes\overline{\textup{bo}}^{\otimes f})\implies\pi_{*}\mathbb{S}_{2}^{\wedge},$
where we may identify $\mathbb{S}_{\text{bo}}^{\wedge}\simeq\mathbb{S}_{2}^{\wedge}$ in the 2-complete setting [Bou79]. There is a complex Betti realization functor
@@
Theorem 2.3.
[ARØ20, Lemma 2.13] If $F$ admits a complex embedding, then there is an equivalence of spectra
@@
Betti realization is symmetric monoidal and exact [HO16], giving an immediate corollary:
Corollary 2.4.
[CQ21, Cor 2.7] Betti realization sends the kq-resolution to the bo-resolution. In particular, there is a multiplicative map of spectral sequences:
$E_{r}^{s,f,*}\to E_{r}^{s,f}.$
@@
Note that, just as in the classical case, the spectrum kq is not flat in the sense of Adams. Thus the $E_{1}$-page does not identify with the cobar complex computing the Hopf algebroid cohomology $\text{Ext}_{\pi^{F}_{**}\textup{kq}\otimes\textup{kq}}(\pi^{F}_{**}\textup{kq},\pi^{F}_{**}\textup{kq})$, and so we must analyze the $E_{1}$-page of the kq-resolution.
Remark 2.6.
There is a complex Betti realization functor in the case of $F=\mathbb{R}$ to which we may apply 2.3. It is important to note that there is an obvious real Betti realization
@@
which is induced by the real analytic topology on a scheme $X\in\text{Sm}_{\mathbb{R}}$ after taking $\mathbb{R}$-points. These functors behave very differently. For example, consider $\text{kq}\in\text{SH}(\mathbb{R})$. It was shown in [Ban25, Proposition 4.1] that $\pi_{*}(\text{Be}^{\mathbb{R}}(\text{kq}))\cong\mathbb{Z}[x]$, where $x\in\pi_{4}(\text{Be}^{\mathbb{R}}(\text{kq}))$. In particular, $\text{Be}^{\mathbb{C}}(\text{kq})\not\simeq\text{Be}^{\mathbb{R}}(\text{kq}).$
@@ 2.2. The dual motivic Steenrod algebra
Let $\mathbb{M}_{2}^{F}$ denote the mod-2 motivic homology of a point. There is an isomorphism due to Voevodsky [Voe03a]:
@@
Theorem 2.7.
[Voe03b, Section 12] The dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{F}=\pi_{**}^{F}(\textup{H}\mathbb{F}_{2}\otimes\textup{H}\mathbb{F}_{2})$ is given by
@@
$(\mathcal{A}\operatorname{/\!/}\mathcal{A}(n))^{\vee}_{F}=\mathbb{M}_{2}^{\mathbb{R}}[\overline{\xi}_{1}^{2^{n}},\overline{\xi}_{2}^{2^{n-1}},\ldots,\overline{\tau}_{n+1},\ldots]/(\overline{\tau}_{i}^{2}=\rho\overline{\tau}_{i+1}+\rho\overline{\tau_{0}}\overline{\xi}_{i+1}+\tau\overline{\xi}_{i+1}).$
@@
An important distinction between $\mathcal{A}^{\vee}_{\mathbb{R}}$ and $\mathcal{A}^{\vee}_{\mathbb{C}}$ is that while the latter is a Hopf algebra, the former is a Hopf algebroid [Rav86, Appendix A]. There is a canonical class $\rho\in\text{K}^{\text{M}}_{1}(F)/2$ representing $-1$, giving rise to a class in $\mathbb{M}_{2}^{F}$. If $-1$ is a square in $F$, then this class is trivial in motivic homology and $\mathcal{A}^{\vee}_{F}$ is an honest Hopf algebra. If $-1$ is not a square, then there is an action of the motivic Steenrod algebra:
@@
Theorem 2.10.
[ARØ20, Remark 2.14] There is an isomorphism of $\mathcal{A}^{\vee}_{F}$-comodule algebras
@@
Proposition 2.11.
[CQ21, Corollary 3.6] The $\mathcal{A}(1)^{\vee}_{F}$-coaction on $\textup{H}_{**}(\textup{kq})$ is determined by the following formulae:
@@
It was shown by Voevodsky [Voe03b] that the motivic Steenrod algebra $\mathcal{A}_{F}$ is free as a module over the cohomology of a point. One basis for this algebra is the motivic Milnor basis [Kyl17]. Since both $\text{Sq}^{1}$ and $\text{Sq}^{2}$ are elements of this basis, the quotient $\mathcal{A}\operatorname{/\!/}\mathcal{A}(1)_{F}\cong\mathcal{A}_{F}\otimes_{\mathcal{A}(1)_{F}}\mathbb{M}_{2}^{F}$ is also free over the cohomology of a point. Consider the following Künneth spectral sequence [DI05]:
$E_{2}=\text{Tor}^{\mathbb{M}_{2}^{F}}(\textup{H}_{**}(\textup{kq}),\textup{H}_{**}(\textup{kq}))\implies\textup{H}_{**}(\textup{kq}\otimes\textup{kq}).$
Since $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))_{F}^{\vee}$ is the $\mathbb{M}_{2}^{F}$-linear dual of a free module, it is also free. Thus the higher Tor vanishes and the spectral sequence collapses with no extensions, giving the result. ∎
@@ 2.3. Motivic Brown–Gitler comodules
and letting $wt(xy)=wt(x)+wt(y)$. This naturally extends to the subalgebras $(\mathcal{A}\operatorname{/\!/}\mathcal{A}(n))_{F}^{\vee}$. The $F$-motivic Brown–Gitler comodule $B_{n}^{F}(k)$ is the $A(n)^{\vee}_{F}$-comodule
@@
One can also define Brown–Gitler comodules using the same weight filtration on the subalgebras $(\mathcal{A}\operatorname{/\!/}\mathcal{E}(n))^{\vee}$, leading to a very different family of comodules. As we will only be working with the case defined above, there should be no ambiguity in the notation.
@@
Definition 2.14.
(compare with [CQ21, Def. 3.10]) The $k$-th integral motivic Brown–Gitler comodule is
@@
We remark that in [CQ21], different notation is used for the Brown–Gitler comodules.
@@
In other words, $B_{0}^{F}(1)$ is presented as the algebraic cell complex in Fig. 1.
$\overline{\xi}_{1}$$\overline{\tau}_{1}$
Figure 1. The motivic Brown–Gitler comodule $B_{0}^{F}(1)$
In fact, we can realize $B_{0}^{F}(1)$ as the mod-2 motivic cohomology of a spectrum [CQ21, Example 3.12]. There is a geometric classifying space [MV99, Section 4] $B\mu_{2}$ of the affine group scheme $\mu_{2}$ of second roots of unity. Take $L_{2}$ to be the simplicial 4-skeleton of $B\mu_{2}$. The inclusion of the $(3,2)$ cell of $L_{2}$ gives a map
@@
In classical topology there exist integral and bo Brown–Gitler spectra [Shi83, GJM86]. These are finite spectra $\textup{H}\mathbb{Z}_{k}$ and $\text{bo}_{k}$, respectively, with the property that
$\text{H}_{*}(\textup{H}\mathbb{Z}_{k})\cong B^{cl}_{0}(k),\quad\text{H}_{*}(\text{bo}_{k})\cong B_{1}^{cl}(k).$
However, while $C_{2}$-equivariant analogues exist for integral Brown–Gitler spectra [LPT24], there are currently no constructions of any motivic Brown–Gitler spectra. One possible construction could come from the motivic lambda algebra of Balderamma–Culver–Quigley [BCQ23] and using the classical work of Goerss–Jones–Mahowald [GJM86] as inspiration. An alternative construction, at least over $\mathbb{C}$, is in using the theory of filtered spectra in [GIKR22]. We will explore the construction of motivic Brown-Gitler spectra in future work.
@@
A key property of the integral motivic Brown–Gitler comodules is that they give rise to a splitting of the homology $\textup{H}_{**}(\textup{kq})$. We refer the reader to [CQ21] for a proof.
Theorem 2.19.
[CQ21, Theorem 3.20] There is an isomorphism of $\mathcal{A}(1)_{F}^{\vee}$-comodules
@@
$\cdots\to\textup{H}_{**}(\mathbb{S})\to\textup{H}_{**}(\textup{kq})\to\textup{H}_{**}(\overline{\textup{kq}})\to H_{*,*-1}(\mathbb{S})\to\cdots.$
We have isomorphisms $\textup{H}_{**}(\mathbb{S})\cong\mathbb{M}_{2}^{F}\cong B_{0}^{F}(0)$ and $\textup{H}_{**}(\textup{kq})\cong\bigoplus_{k\geq 0}\Sigma^{4k,2k}B_{0}^{F}(k)$. The unit map induces an inclusion of the bottom summand $B_{0}^{F}(0)\hookrightarrow\bigoplus_{k\geq 0}B_{0}^{F}(k)$, so by the long exact sequence, $\textup{H}_{**}(\overline{\textup{kq}})$ is trivial in those degrees. The long exact sequence also gives an isomorphism between $\textup{H}_{**}(\textup{kq})$ and $\textup{H}_{**}(\overline{\textup{kq}})$ outside of these degrees since $\textup{H}_{**}(\mathbb{S})$ vanishes, giving the result. ∎
We close this section by recalling useful short exact sequences of $\mathcal{A}(1)^{\vee}_{F}$-comodules relating motivic Brown–Gitler comodules. We refer the reader to [CQ21] for a proof.
Proposition 2.21.
[CQ21, Lemma 3.21] There are short exact sequences of $\mathcal{A}(1)^{\vee}_{F}$-comodules:
@@
$0\to\Sigma^{4k,2k}B_{0}^{F}(k)\otimes_{\mathbb{M}_{2}^{F}}B_{0}^{F}(1)\to B_{0}^{F}(2k+1)\to B_{1}^{F}(k-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{F}^{\vee}\to 0$
@@
The short exact sequences in Proposition 2.21 are motivic analogues of classical short exact sequences of Brown–Gitler comodules. These arise by applying mod-2 homology to cofiber sequences relating integral and bo Brown–Gitler spectra [Pea14]. One would imagine the same to be true for the conjectural motivic Brown–Gitler spectra.

### v2

@@ 1.4. Organization
Part I of this paper presents background and recalls the work of Culver–Quigley in [17]. In Section 2, we recall relevant facts about the stable motivic homotopy category $\text{SH}(F)$, the kq-resolution, and motivic Brown–Gitler comodules. In Section 3, we prove a novel result about symplectic K-theory. In Section 4, we outline a process for computing the ring of cooperations $\pi_{**}^{F}(\text{kq}\otimes\text{kq})$. Much of the material of these sections is contained [17, Section 3], but we phrase things in slightly more generality, which will be useful to a broader audience. In Section 5, we review the computation of $\pi_{**}^{\mathbb{C}}(\text{kq}\otimes\text{kq})$ as a warm up for the main contents of our work.
Part II of this paper contains the computation of the ring of cooperations $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$ and applications. In Section 6, we describe the algebra $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})$ and some modules which will frequently appear. In Section 7, we compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(\mathbb{M}_{2}^{\mathbb{R}},B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ by a series of algebraic Atiyah–Hirzebruch spectral sequences. In Section 8, we compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\text{kq}\otimes\text{kq})$ and apply our results to describe the $\textup{E}_{1}$-page of the kq-resolution.
@@ 1.5. Notation and Conventions
  • KQ denotes the motivic spectrum representing Hermitian K-theory, kq denotes its very effective cover, and ksp denotes the very effective cover of $\Sigma^{4,2}\text{KQ}$.
  • $\text{Ext}^{***}_{B}(\mathbb{M}_{2}^{F},X)$ will be abbreviated by $\text{Ext}^{***}_{B}(X)$.
  • We grade Ext groups as $(s,f,w)$, where $s$ is the stem, $f$ is the Adams filtration, and $w$ is the motivic weight. We also let cw=swcw=s-w denote the coweight, sometimes referred to in the literature as the Milnor-Witt degree.
@@
  • For E a motivic ring spectrum, we let $\text{E-}\textbf{mASS}^{F}(X)$ denote the motivic Adams spectral sequence based on E and converging to $\pi_{**}^{F}(X_{\text{E}}^{\wedge})$. When $\text{E}=\text{H}\mathbb{F}_{2}$, we denote this spectral sequence by $\textbf{mASS}^{F}(X).$
@@ 1.6. Acknowledgments
The work presented here constitutes a part of the author’s thesis. The author thanks their advisors, Kyle Ormsby and John Palmieri, for their indispensable wisdom and valuable advice. The author also thanks J.D. Quigley for his generosity in sharing insight into this problem, and thanks Madeline Borowski, Sarah Petersen, Jay Reiter, and Alex Waugh for helpful discussions throughout this project. Finally, the author thanks Noah Wisdom for providing valuable feedback on an early draft, MJ Lenderman for inspiration [37], and an anonymous referee for helpful suggestions which clarified many technical points.
@@ 2.1. The kq-resolution
Originally constructed in [27] over fields with 2 invertible, and more recently, over more general base schemes in [15], there is an $(8,4)$-periodic motivic $\mathbb{E}_{\infty}$-ring spectrum $\textup{KQ}\in\text{SH}(F)$ known as the Hermitian K-theory spectrum. This spectrum represents Hermitian K-theory in that there is an isomorphism [15, Theorem 8.18]
$\pi^{F}_{s,w}(\textup{KQ})=\pi_{s-2w}(\textup{GW}^{\text{s}}(F)),$
where $\text{GW}^{\text{s}}(F)$ is the ordinary spectrum representing the symmetric Hermitian K-theory of $F$. In particular, $\pi_{8n,4n}^{F}\textup{KQ}=\textup{GW}(F)$, the Grothendieck–Witt ring of quadratic forms over $F$, and the unit map $\mathbb{S}\to\text{KQ}$ induces Morel’s isomorphism [50]
$\pi^{F}_{-n,-n}\mathbb{S}\cong\textup{K}^{\textup{MW}}_{n}(F),$
where $\text{K}^{\text{MW}}_{n}(F)$ denotes the $n^{th}$ Milnor K-theory of $F$.
Let kq denote the very effective cover of KQ [1] (see Section 3 for more details on the very effective filtration), which we will call the very effective Hermitian K-theory spectrum. There is a canonical Adams tower associated to the unit map $\mathbb{S}\to\textup{kq}$:
@@
Applying $\pi^{F}_{**}(-)$ yields the motivic Adams spectral sequence kq-$\textbf{mASS}^{F}(\mathbb{S})$ known as the kq-resolution. Its properties were first studied by Culver–Quigley [17]. We remark that we are working in the 2-complete category.
Theorem 2.1 ([17, Thm 2.1]).
The kq-resolution is a strongly convergent spectral sequence of the form
$\textup{E}_{1}^{s,f,w}=\pi^{F}_{s+f,w}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes f})\implies\pi^{F}_{s,w}\mathbb{S}_{2}^{\wedge}.$
@@
We can also look at the kq-resolution in the integral, uncompleted setting. The unit map $\mathbb{S}\to\textup{kq}$ gives an isomorphism
$\Pi_{0}^{F}(\textup{kq})\cong\Pi_{0}^{F}(\mathbb{S}),$
where $\Pi_{n}^{F}(X):=\bigoplus_{k\in\mathbb{Z}}\pi_{n+k,k}^{F}(X)$ denotes the $k$-th coweight stem (also known as the $k$-th Milnor–Witt stem). This implies that $\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}$ is $n$-connected in Morel’s homotopy $t$-structure. In fact, this is enough to show that $\mathbb{S}_{\textup{kq}}^{\wedge}\simeq\mathbb{S}$.
However, we will compute the $\textup{E}_{1}$-page of the kq-resolution by a motivic Adams spectral sequence based on the Eilenberg–Mac Lane spectrum $\textup{H}\mathbb{F}_{2}$. More precisely, we study te $\textbf{{mASS}}^{F}(\text{kq}\otimes\text{kq})$. This spectral sequence computes the homotopy groups of the $(2,\eta)$-completion of $\textup{kq}\otimes\textup{kq}$. When $F$ is a field of finite virtual 2-cohomological dimension, then this coincides with the 2-completion [28, Theorem 1]. These conventions force us to work in the 2-complete category (see also [17, Section 2.2]).
Let bo denote the connective cover of the classical real K-theory spectrum KO. The kq-resolution is the motivic analogue of the bo-resolution in the following way. The $\textup{E}_{1}$-page of the bo-resolution is of the form
$\textup{E}_{1}^{s,f}=\pi_{s+f}(\textup{bo}\otimes\overline{\textup{bo}}^{\otimes f})\implies\pi_{s}(\mathbb{S}_{2}^{\wedge}),$
where we may identify $\mathbb{S}_{\text{bo}}^{\wedge}\simeq\mathbb{S}_{2}^{\wedge}$ in the 2-complete setting [13]. There is a complex Betti realization functor
@@
Theorem 2.2 ([1, Lemma 2.13]).
If $F$ admits a complex embedding, then there is an equivalence of spectra
@@
Betti realization is symmetric monoidal and exact [25], giving an immediate corollary:
Corollary 2.3 ([17, Cor 2.7]).
Betti realization sends the kq-resolution to the bo-resolution. In particular, there is a multiplicative map of spectral sequences:
$\textup{E}_{r}^{s,f,*}\to\textup{E}_{r}^{s,f}.$
Just as in the classical case, the spectrum kq is not flat in the sense of Adams. Thus the $\textup{E}_{1}$-page does not compute the Hopf algebroid cohomology $\text{Ext}^{***}_{\pi^{F}_{**}(\textup{kq}\otimes\textup{kq})}(\pi^{F}_{**}(\textup{kq}),\pi^{F}_{**}(\textup{kq}))$, and so we must analyze the $\textup{E}_{1}$-page of the kq-resolution.
@@
There is a complex Betti realization functor in the case of $F=\mathbb{R}$ to which we may apply 2.2. It is important to note that there is an obvious real Betti realization
@@
which is induced by the real analytic topology on a scheme $X\in\text{Sm}_{\mathbb{R}}$ after taking $\mathbb{R}$-points. These functors behave very differently. For example, consider $\text{kq}\in\text{SH}(\mathbb{R})$. It was shown in [8, Proposition 4.1] that $\pi_{*}(\text{Be}^{\mathbb{R}}(\text{kq}))\cong\mathbb{Z}[x]$, where $x\in\pi_{4}(\text{Be}^{\mathbb{R}}(\text{kq}))$. In particular, $\text{Be}^{\mathbb{C}}(\text{kq})\not\simeq\text{Be}^{\mathbb{R}}(\text{kq}).$
@@ 2.2. The dual motivic Steenrod algebra
Let $\mathbb{M}_{2}^{F}$ denote the mod-2 motivic homology of a point. There is an isomorphism due to Voevodsky [55]:
@@
Theorem 2.4 ([56, Section 12]).
The dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{F}=\pi_{**}^{F}(\textup{H}\mathbb{F}_{2}\otimes\textup{H}\mathbb{F}_{2})$ is given by
@@
$(\mathcal{A}\operatorname{/\!/}\mathcal{A}(n))^{\vee}_{F}=\mathbb{M}_{2}^{F}[\overline{\xi}_{1}^{2^{n}},\overline{\xi}_{2}^{2^{n-1}},\ldots,\overline{\tau}_{n+1},\ldots]/(\overline{\tau}_{i}^{2}=\rho\overline{\tau}_{i+1}+\rho\overline{\tau_{0}}\overline{\xi}_{i+1}+\tau\overline{\xi}_{i+1}).$
@@
An important distinction between $\mathcal{A}^{\vee}_{\mathbb{R}}$ and $\mathcal{A}^{\vee}_{\mathbb{C}}$ is that while the latter is a Hopf algebra, the former is a Hopf algebroid [49, Appendix A]. There is a canonical class $\rho\in\text{K}^{\text{M}}_{1}(F)/2$ representing 1-1, giving rise to a class in $\mathbb{M}_{2}^{F}$. If 1-1 is a square in $F$, then this class is trivial in motivic homology and $\mathcal{A}^{\vee}_{F}$ is an honest Hopf algebra. If 1-1 is not a square, then there is an action of the motivic Steenrod algebra:
@@
Theorem 2.5 ([1, Remark 2.14]).
There is an isomorphism of $\mathcal{A}^{\vee}_{F}$-comodule algebras
@@
Proposition 2.6 ([17, Corollary 3.6]).
The $\mathcal{A}(1)^{\vee}_{F}$-coaction on $\textup{H}_{**}(\textup{kq})$ is determined by the following formulae:
@@
It was shown by Voevodsky [56] that the motivic Steenrod algebra $\mathcal{A}_{F}$ is free as a module over the cohomology of a point. One basis for this algebra is the motivic Milnor basis [56, 35]. Since both $\text{Sq}^{1}$ and $\text{Sq}^{2}$ are elements of this basis, the quotient $\mathcal{A}\operatorname{/\!/}\mathcal{A}(1)_{F}\cong\mathcal{A}_{F}\otimes_{\mathcal{A}(1)_{F}}\mathbb{M}_{2}^{F}$ is also free over the cohomology of a point. Consider the following Künneth spectral sequence [18]:
$\textup{E}_{2}=\text{Tor}^{\mathbb{M}_{2}^{F}}(\textup{H}_{**}(\textup{kq}),\textup{H}_{**}(\textup{kq}))\implies\textup{H}_{**}(\textup{kq}\otimes\textup{kq}).$
Since $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))_{F}^{\vee}$ is the $\mathbb{M}_{2}^{F}$-linear dual of a finitely-generated free module, it is also free. Thus the higher Tor vanishes and the spectral sequence collapses with no extensions, giving the result. ∎
@@ 2.3. Motivic Brown–Gitler comodules
and letting $wt(xy)=wt(x)+wt(y)$. This naturally extends to the subalgebras $(\mathcal{A}\operatorname{/\!/}\mathcal{A}(n))_{F}^{\vee}$. The $F$-motivic Brown–Gitler comodule $B_{n}^{F}(k)$ is the $\mathcal{A}(n)^{\vee}_{F}$-comodule
@@
One can also define Brown–Gitler comodules using the same weight filtration on the subalgebras $(\mathcal{A}\operatorname{/\!/}\mathcal{E}(n))_{F}^{\vee}$, leading to a different family of comodules. These families of Brown–Gitler comodules were used by the author, Petersen and Tatum in [45], particularly in the $n=0$ case. As we will only be working with the case defined above, and since $\mathcal{A}(0)\cong\mathcal{E}(0)$, there should be no ambiguity in the notation.
@@
Definition 2.6 (compare with [17, Def. 3.10]).
The $k$-th integral motivic Brown–Gitler comodule is
@@
We remark that in [17], different notation is used for the Brown–Gitler comodules.
@@
In other words, $B_{0}^{F}(1)$ is presented as the algebraic cell complex in Figure 2.1.
$h_{1}$$h_{0}$
Figure 2.1. The motivic Brown–Gitler comodule $B_{0}^{F}(1)$
In fact, we can realize $B_{0}^{F}(1)$ as the mod-2 motivic cohomology of a spectrum [17, Example 3.12]. There is a geometric classifying space [42, Section 4] $B\mu_{2}$ of the affine group scheme $\mu_{2}$ of second roots of unity. Take $L_{2}$ to be the simplicial 4-skeleton of $B\mu_{2}$. The inclusion of the $(3,2)$ cell of $L_{2}$ gives a map
@@
In classical topology there exist integral and bo Brown–Gitler spectra [52, 23]. These are finite spectra $\textup{H}\mathbb{Z}_{k}$ and $\text{bo}_{k}$, respectively, with the property that
$\text{H}_{*}(\textup{H}\mathbb{Z}_{k})\cong B^{\text{cl}}_{0}(k),\quad\text{H}_{*}(\text{bo}_{k})\cong B_{1}^{\text{cl}}(k).$
However, while $C_{2}$-equivariant analogues exist for integral Brown–Gitler spectra [39], there are currently no constructions of any motivic Brown–Gitler spectra. One possible construction could come from the motivic lambda algebra of Balderamma–Culver–Quigley [7] and using the classical work of Goerss–Jones–Mahowald [23] as inspiration. An alternative construction, at least over $\mathbb{C}$, is in using the theory of filtered spectra in [21]. We will explore the construction of motivic Brown–Gitler spectra in future work.
@@
A key property of the integral motivic Brown–Gitler comodules is that they give rise to a splitting of the homology $\textup{H}_{**}(\textup{kq})$. We refer the reader to [17] for a proof.
Theorem 2.9 ([17, Theorem 3.20]).
There is an isomorphism of $\mathcal{A}(1)_{F}^{\vee}$-comodules
@@
$\cdots\to\textup{H}_{**}(\mathbb{S})\to\textup{H}_{**}(\textup{kq})\to\textup{H}_{**}(\overline{\textup{kq}})\to\text{H}_{*-1,*}(\mathbb{S})\to\cdots.$
We have isomorphisms
$\textup{H}_{**}(\mathbb{S})\cong\mathbb{M}_{2}^{F}\cong B_{0}^{F}(0)\quad\text{and}\quad\textup{H}_{**}(\textup{kq})\cong\bigoplus_{k\geq 0}\Sigma^{4k,2k}B_{0}^{F}(k).$
The unit map $\mathbb{S}\to\text{kq}$ induces an inclusion of the bottom summand $B_{0}^{F}(0)\hookrightarrow\bigoplus_{k\geq 0}B_{0}^{F}(k)$, so by the long exact sequence, $\textup{H}_{**}(\overline{\textup{kq}})$ is trivial in those degrees. The long exact sequence also gives an isomorphism between $\textup{H}_{**}(\textup{kq})$ and $\textup{H}_{**}(\overline{\textup{kq}})$ outside of these degrees since $\textup{H}_{**}(\mathbb{S})$ vanishes, giving the result. ∎
We close this section by recalling useful short exact sequences of $\mathcal{A}(1)^{\vee}_{F}$-comodules relating motivic Brown–Gitler comodules. We refer the reader to [17] for a proof.
Proposition 2.11 ([17, Lemma 3.21]).
There are short exact sequences of $\mathcal{A}(1)^{\vee}_{F}$-comodules:
@@
$0\to\Sigma^{4k,2k}B_{0}^{F}(k)\otimes_{\mathbb{M}_{2}^{F}}B_{0}^{F}(1)\to B_{0}^{F}(2k+1)\to B_{1}^{F}(k-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{F}^{\vee}\to 0.$
@@
The short exact sequences in Proposition 2.11 are motivic analogues of classical short exact sequences of Brown–Gitler comodules. These arise by applying mod-2 homology to cofiber sequences relating integral and bo Brown–Gitler spectra [47]. One would imagine the same to be true for the conjectural motivic Brown–Gitler spectra.